OpenStudy (anonymous):

A 0.100μg speck of dust is accelerated from rest to a speed of 0.900 c by a constant 1.00×10^6N force.

OpenStudy (anonymous):

If the nonrelativistic form of Newton's second law F=ma is used, how far does the object travel to reach its final speed?

OpenStudy (anonymous):

@theEric Hey, do you know how to do this?

OpenStudy (theeric):

Hi! Well, if your not taking special relativity into account, you can use fundamental kinematic equations! You know the mass, force, and initial and final velocity. From the force and mass, you can get the acceleration. You can skip to a derived equation with acceleration, initial and final velocity, and displacement to answer this. To go on to applying special relativity, I would need a refresher.

OpenStudy (anonymous):

This involves some sort of relativity but I have no idea what, I see you can get acceleration but how would you go on to get the distance travelled in its final speed?

OpenStudy (theeric):

\(v^2=u^2+2a\Delta x\) I forget how to derive that :P

OpenStudy (anonymous):

Ah so, \[\Delta x = \frac{ v^2 }{ 2a } \implies \frac{ 0.900c }{ 2(9.23*10^{12}m/s^2) }\]

OpenStudy (anonymous):

Sorry the acceleration is wrong, wrong numbers xD

OpenStudy (anonymous):

a = 1*10^13 m/s^2

OpenStudy (anonymous):

Oh and that should be squared...(0.900c)^2 bleh

OpenStudy (theeric):

I'll agree after you also square the numerator to make - oh, you got it :)

OpenStudy (theeric):

Yep. That looks good for classical physics as far as I know!

OpenStudy (anonymous):

\[0.100 \mu g = 0.100*10^{-6} kg?\] right because I think I did something wrong, not sure what..mhm

OpenStudy (theeric):

Ah, that's the issue! Sorry I didn't catch it either! It's just grams, there. So it's \(10^{-3}\) kilograms.

OpenStudy (theeric):

uh, that's the order, not the number.

OpenStudy (anonymous):

micrograms isn't it haha? I need to go look at the metric system I thought it was -6

OpenStudy (theeric):

\(0.1\times10^{-3}=10^{-4}\)

OpenStudy (theeric):

It is! \(10^{-6}\) grams, though! And I messed up!

OpenStudy (theeric):

\(0.1\times10^{-9}\) grams kg g mg \(\mu\)g

OpenStudy (anonymous):

oh yeah it's something like 10^-10 then lol

OpenStudy (anonymous):

One sec, let me use my name to make sure!

OpenStudy (theeric):

Haha, good plan.

OpenStudy (anonymous):

I thought it was ^-10 kg

OpenStudy (anonymous):

Yup there we go haha

OpenStudy (theeric):

Haha, yep! Your name agrees.

OpenStudy (anonymous):

Alright, there's a second part mind helping me with it, I'm not really sure what K means here...Using the correct relativistic treatment (K=(y-1)mc^2) , how far does the object travel to reach its final speed? y=gamma

OpenStudy (anonymous):

Or is K suppose to be kinetic energy

OpenStudy (anonymous):

\[\gamma = \frac{ 1 }{ \sqrt{1-v^2/c^2} }\] I guess to find gamma we could use this, but not sure about K and the distance.

OpenStudy (theeric):

I think that might be relativistic kinetic energy!

OpenStudy (theeric):

And \(\gamma\) is like the Lorentz transformation or something, right?

OpenStudy (anonymous):

Yeah so wait a sec, we have this then \[K = \frac{ mc^2 }{ \sqrt{1-v^2/c^2} }-mc^2 = (\gamma - 1)mc^2\]

OpenStudy (anonymous):

So I can get gamma and figure out K but then for distance?

OpenStudy (theeric):

Good question...

OpenStudy (theeric):

The \(\gamma\) will change with the velocity...

OpenStudy (theeric):

Can we use work at all somehow? The force is doing work, right? But I don't know how to deal with special relativity!

OpenStudy (anonymous):

I'm sure we can since \[K = W = \int\limits_{0}^{v} \frac{ mv_x dv_x}{ (1-v_x ^2/c^2)^{3/2} }\]

OpenStudy (anonymous):

That's how you would derive the expression above.

OpenStudy (anonymous):

\[W = Fd ?\] can we just use that? Or do we have to integrate

OpenStudy (anonymous):

W = K so the expressions for K/F = d

OpenStudy (anonymous):

This is a pretty cool question as far as comparing classical and relativistic aspects will be neat to see the final answer on which one has a greater distance.

OpenStudy (theeric):

Energy is conserved, even in special relativity, right? So to have a gain in kinetic energy that you calculate with your equation, that much work has to be done. It's done by that given force. So I think it is that simple! :)

OpenStudy (theeric):

I agree! :)

OpenStudy (theeric):

Since we're looking at two points only, I don't think integration is necessary :)

OpenStudy (anonymous):

Yup, it was right, and relativity give a large number, neat...

OpenStudy (theeric):

Cool! What was the order of the large number?

OpenStudy (theeric):

\(10^?\)

OpenStudy (anonymous):

Not by much actually like I was using different numbers as this one was straight from the book but classically I got around 4m and with relativity 14m, still neat I guess, I'm assuming relativity is more accurate?